OpenStudy (anonymous):

If a hotel room normally cost $200 per night and you are getting it at $69 per night, then what would be the discount percentage that you are saving? (round to the nearest tenth percent.)

OpenStudy (anonymous):

discount percent = difference in price/original price discounted percent = (200 - 69)/200 = 131/200 = 65.5% discount

OpenStudy (anonymous):

oh ok. I was doing it backwards because I had 34.5%. But I knew that was not right. I get it now thank you very much. ;-)